OpenStudy (anonymous):

Let h(x)=2x+5 and g(x)=8−3x What is the value of h(g(1.5))?

OpenStudy (loser66):

Where are you stuck?

OpenStudy (anonymous):

the last equation how to solve it

OpenStudy (loser66):

if g(x) = 8-3x, then g(1.5) =??

OpenStudy (anonymous):

3.5

OpenStudy (loser66):

YYYYYYYYYYYes

OpenStudy (loser66):

Now, if h (x) =2x+5, then h(3.5) =?

OpenStudy (anonymous):

12

OpenStudy (loser66):

you got it.

OpenStudy (anonymous):

is that really all it was

OpenStudy (loser66):

yes, dat sit

OpenStudy (loser66):